# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ B.2 Omitted Proofs of Section˜4
$\sqrt{n}\,(\widehat{\tau}_{\textup{LDM}}-\tau)\quad\text{and}\quad\sqrt{n}\,(\widehat{\tau}^{*}_{\textup{LDM}}-\tau)$
have the same asymptotic distribution. It remains to analyze $\widehat{\tau}^{*}_{\textup{LDM}}$.
Now note that since $\boldsymbol{\beta}_{\lambda}^{*}$ is a fixed vector, by (13), the limiting variance of $\widehat{\tau}^{*}_{\textup{LDM}}$ is given by
$\lim_{n\to\infty}\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{*}_{\lambda}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2}.$
Because complete random assignment corresponds to sampling $n_{T}$ treated units without replacement from a finite population of fixed potential outcomes and covariates, and because $n_{T}/n$ and $n_{C}/n$ are bounded away from 0 and $1$ by Assumption˜3, Hájek’s finite-population central limit theorem for sampling without replacement (hajek1960limiting) applies to $\sqrt{n}\,(\widehat{\tau}^{*}_{\textup{LDM}}-\tau)$, under our bounded fourth-moment conditions (Assumptions˜1 and 2). In particular,
$\sqrt{n}\,(\widehat{\tau}^{*}_{\textup{LDM}}-\tau)\;\xrightarrow{d}\;\mathcal{N}\!\Big(0,\;\lim_{n\to\infty}\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{*}_{\lambda}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2}\Big).$
Finally, since $\widehat{\tau}_{\textup{LDM}}-\widehat{\tau}^{*}_{\textup{LDM}}\;\xrightarrow{p}\;0$, Slutsky’s theorem implies that the same asymptotic normal limit holds for $\sqrt{n}(\widehat{\tau}_{\textup{LDM}}-\tau)$. ∎

### v2

@@ Appendix C Auxilary Lemmas
$\mathbb{E}\left[\|S\|_{2}^{2}\right]$ $=\alpha^{2}\,\mathbb{E}\left[\left\|\sum_{i=1}^{n}(d_{i}-p_{T}){\bm{y}}_{i}\right\|_{2}^{2}\right]$
$=\alpha^{2}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbb{E}\big[(d_{i}-p_{T})(d_{j}-p_{T})\big]\;{\bm{y}}_{i}^{\top}{\bm{y}}_{j}$
$=\alpha^{2}\left(\sum_{i=1}^{n}p_{T}(1-p_{T})\|{\bm{y}}_{i}\|_{2}^{2}+\sum_{i\neq j}\left(-\frac{p_{T}(1-p_{T})}{n-1}\right){\bm{y}}_{i}^{\top}{\bm{y}}_{j}\right).$
Since $\sum_{i}{\bm{y}}_{i}=\boldsymbol{0}$,
$\sum_{i\neq j}{\bm{y}}_{i}^{\top}{\bm{y}}_{j}=\left\|\sum_{i=1}^{n}{\bm{y}}_{i}\right\|_{2}^{2}-\sum_{i=1}^{n}\|{\bm{y}}_{i}\|_{2}^{2}=-\sum_{i=1}^{n}\|{\bm{y}}_{i}\|_{2}^{2}.$
Therefore
$\mathbb{E}\left[\|S\|_{2}^{2}\right]=\alpha^{2}\,p_{T}(1-p_{T})\left(1+\frac{1}{n-1}\right)\sum_{i=1}^{n}\|{\bm{y}}_{i}\|_{2}^{2}=\alpha^{2}\,p_{T}(1-p_{T})\frac{n}{n-1}\sum_{i=1}^{n}\|{\bm{x}}_{i}-\overline{{\bm{x}}}\|_{2}^{2}.$
Finally, $\sum_{i=1}^{n}\|{\bm{x}}_{i}-\overline{{\bm{x}}}\|_{2}^{2}\leq\sum_{i=1}^{n}\|{\bm{x}}_{i}\|_{2}^{2}\leq Cn$, hence
$\mathbb{E}\|S\|_{2}^{2}\leq\alpha^{2}\,p_{T}(1-p_{T})\frac{n}{n-1}\,Cn.$
Since $\alpha=\frac{n}{n_{T}n_{C}}$ and $p_{T}(1-p_{T})=\frac{n_{T}n_{C}}{n^{2}}$, this simplifies to
$\mathbb{E}\|S\|_{2}^{2}\leq\frac{C}{n-1}\cdot\frac{n^{2}}{n_{T}n_{C}}=O(\frac{n}{n_{T}n_{C}}).$
Then by Markov’s inequality, for any $M>0$,
$\mathbb{P}\!\left(\|S\|_{2}\geq M\sqrt{\frac{n}{n_{T}n_{C}}}\right)=\mathbb{P}\!\left(\|S\|_{2}^{2}\geq\frac{M^{2}n}{n_{T}n_{C}}\right)\leq\frac{\mathbb{E}\|S\|_{2}^{2}}{\frac{M^{2}n}{n_{T}n_{C}}}=O(\frac{1}{M^{2}}).$
Given any $\varepsilon>0$, we can choose $M:=M(\varepsilon)$ large enough so that the right-hand-side above is smaller than $\varepsilon$. Therefore $\|S\|_{2}=O_{p}\!\left(\sqrt{\frac{n}{n_{T}n_{C}}}\right)$, and under Assumption˜2, since $n_{T}/n$ and $n_{C}/n$ are bounded away from zero and one, $\|S\|_{2}=O_{p}(n^{-1/2})$. ∎
Appendix D Efficiency of LOORA-DM
In this section, we study the asymptotic efficiency of the LOORA-DM estimator. Since the estimator of Lin (2013) with interacted terms is asymptotically efficient among linearly adjusted estimators (see Cytrynbaum (2024)), it suffices to show that LOORA-DM is asymptotically as efficient as the estimator of Lin (2013).
The estimator of Lin (2013), denoted by $\widehat{\tau}_{\textup{Lin}}$, is obtained by regressing the observed outcome vector on $\bm{d}$}, $\bm{X}$}, and $\boldsymbol{D}({\bm{X}}-\overline{{\bm{X}}})$, with an intercept, where $\boldsymbol{D}$ denotes the diagonal matrix formed from $\bm{d}$}. Under complete random assignment (with $n_{T}$ units assigned to treatment and $n_{T}/n\to p_{T}$) and suitable regularity conditions, the asymptotic variance of the estimator is given by
$\lim_{n\to\infty}{\rm Var}(\sqrt{n}\widehat{\tau}_{\textup{Lin}})$ $=\frac{1-p_{T}}{p_{T}}\lim_{n\to\infty}\sigma_{n}^{2}({\bm{y}}^{(1)})+\frac{p_{T}}{1-p_{T}}\lim_{n\to\infty}\sigma_{n}^{2}({\bm{y}}^{(0)})\ +2\lim_{n\to\infty}\sigma_{n}({\bm{y}}^{(1)},{\bm{y}}^{(0)}),$
where
$\sigma_{n}^{2}({\bm{y}}^{(1)})=\frac{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{(1)}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\right\|_{2}^{2}}{n},~\boldsymbol{\beta}^{(1)}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\right\|_{2}^{2},$
$\sigma_{n}^{2}({\bm{y}}^{(0)})=\frac{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{(0)}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\right\|_{2}^{2}}{n},~\boldsymbol{\beta}^{(0)}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\right\|_{2}^{2},$
$\sigma_{n}({\bm{y}}^{(1)},{\bm{y}}^{(0)})=\frac{({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)}-({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{(1)})^{\top}({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)}-({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{(0)})}{n}.$
Lin (2013) considers only the case of complete random assignment and a regression adjustment without regularization (i.e, $\lambda=0$). There are also other differences between our setting and that of Lin (2013). For example, we do not assume that $n_{T}/n$ or the variance of $\sqrt{n}\widehat{\tau}_{\textup{LDM}}$ converges; that is, these quantities may fluctuate with $n$.
[…]
$\sigma_{n}^{2}(\widetilde{{\boldsymbol{\mu}}})=\frac{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}^{(\widetilde{{\boldsymbol{\mu}}})}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}^{2}}{n},~~\boldsymbol{\beta}^{(\widetilde{{\boldsymbol{\mu}}})}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{k}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}^{2}.$
Simple algebra yields
[…]
Therefore LOORA-DM is as efficient as the estimator of Lin (2013).